# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ IV-C3 Impact of V2X on CP
CP demands real-time, high-quality data, often involving high-resolution inputs from sensors such as LiDAR and radar, which places significant data rate requirements on communication systems. Achieving high data rates typically requires large bandwidth, stable transmission channels, and precise channel detection and estimation. DSRC operates within the 5.9 GHz frequency band with a total bandwidth of 75 MHz, divided into seven 10 MHz channels ($5.850-5.925$ GHz). However, due to high vehicle mobility, dynamic environmental changes, and multipath delay spread, DSRC’s communication channels are highly time- and frequency-selective [172]. Studies estimate that up to 50% of the coherence bandwidth is in the order of 1 MHz, with coherence times as short as 0.2 ms [172], limiting reliable transmission times and increasing dropout rates. Similar constraints affect C-V2X, which operates on a narrower 30 MHz spectrum ($5.895-5.925$ GHz) [173]. The PC5 mode of C-V2X offers an advantage by utilizing a blind hybrid automatic repeat request (HARQ) technique that combines turbo coding (providing better performance than the convolutional coding used in DSRC) with automatic repeat request error control and soft-combining [174].
Addressing these challenges requires improved network planning and advanced channel detection and estimation techniques. For DSRC, [175] proposes an analytical model to assess key metrics such as Channel Busy Ratio and Packet Delivery Ratio over distance, facilitating optimized network planning. Additionally, [176] enhances IEEE 802.11p through an OFDM-IM system that strategically positions pilot signals and leverages inactive channels to increase detection accuracy, while [177] introduces a low-complexity, high-speed deep neural network-augmented least square algorithm, effectively reducing model size and latency while improving accuracy under varied noise levels and channel conditions. For C-V2X, [178] proposes a balanced resource allocation strategy that evenly distributes resources across subframes, decreasing packet loss and enhancing data throughput. Similarly, [179] suggests a pilot design with zero-correlation zone sequences and a time-domain channel estimation approach to mitigate co-channel interference in C-V2X systems.
@@
To address congestion in DSRC, [181] proposes adjustments to vehicle settings based on information from supplementary channels, enabling lead vehicles to monitor hidden nodes in a group. Furthermore, [182] presents a congestion-aware forwarding scheme that optimizes multi-hop forwarding reliability and reduces latency under heavy network load. In C-V2X, [179] introduces a rate control mechanism that allows power control to actively contribute to congestion management, while [183] suggests a DQN-based adaptive congestion control algorithm, where a central controller trains and distributes the policy, improving QoS compliance over ETSI’s Distributed Congestion Control in dense traffic scenarios. For synchronization, [184] offers a decentralized clock synchronization protocol to enhance DSRC’s resilience against GPS spoofing and improve reliability when GNSS signals are unavailable. Meanwhile, [185] proposes a synchronization signal design that uses generalized reciprocal sequences for mmWave C-V2X. This approach enhances OFDM-based V2V synchronization by achieving high detection accuracy, minimizing interference, and reducing computational complexity.
@@ IV-D Cooperative Information Alignment
Cooperative information alignment is crucial for CP systems in autonomous driving to ensure that shared data from different sources are accurately represented within a common coordinate system [186]. This process is essential for fusing information from multiple agents to achieve a unified understanding of the environment.
@@ IV-D2 Alignment for Pose Errors
TABLE IV: Summary of the Works of Cooperative Information Alignment for Pose Errors.
@@
DMGM [199] 2023 V2V Camera CARLA & SUMO Early -- $\surd$
@@
Many efforts have been undertaken to address pose errors, as summarized in Table IV. A widely adopted approach involves statistical learning models. For instance, the method presented in [188] corrected relative locations between two CAVs by matching key points in their perception results with the Random Sample Consensus (RANSAC) algorithm [200]. While effective, this approach required the setting of problem-specific thresholds. Based on V2VNet [140], a consistency module [189] was introduced to enhance relative pose estimations by structuring the global coherence of absolute poses as a Markov random field and incorporating Bayesian reweighting. Unfortunately, this model relies on ground-truth poses during the training phase, which may be limited in practical scenarios. CoAlign [190] proposed an Agent-Object Pose Graph Optimization method to improve pose consistency using a pose graph and optimization techniques. RoCo [191] modelled the pose correction problem as an object-matching task that reliably associates common objects detected by different agents, iteratively adjusting the agent poses to minimize alignment errors of the associated objects. These two methods effectively mitigated relative pose errors among agents without requiring precise supervision during training.
@@
Point cloud registration methods typically focus on refining the Iterative Closest Point (ICP) algorithm [201], widely used for aligning 3D point clouds. In [192], an ICP-based matching algorithm aligned two point sets, one representing 3D bounding boxes detected by the Ego and the other by the sender, with the relative transformation formulated as an optimal transport problem [202] to minimize transportation costs between source and target points. V2X-PC [136] aligned point clusters from spatial and temporal dimensions and proposed parameter-free solutions that can adapt to different noisy levels without finetuning. CoBEVGlue [193] introduced a spatial alignment module, BEVGlue, which estimates relative poses between agents by matching co-visible objects across agents without relying on an external localization system. Unlike traditional point cloud registration methods, BEVGlue only requires the transmission of bounding boxes and tracking IDs, reducing communication bandwidth while maintaining high-quality matching by checking the spatial relationship between matched nodes.
@@
High-definition maps represent a valuable resource for achieving precise spatial alignment, primarily due to their ability to provide accurate self-vehicle localization. In this regard, a map-based CP framework called the “map container” was introduced in [194], which establishes benchmarks for spatiotemporal transformations by automatically mapping multi-agent perceptual information into the map coordinate space using a map-matching algorithm. Nevertheless, this algorithm relies on the support of monocular cameras and GNSS/IMU devices. FreeAlign [195] leveraged the invariant geometric structure composed of objects commonly perceived and observed by the individual agents to estimate the relative poses without relying on expensive high-end global localization.
@@
Multi-scale feature interaction can effectively reduce feature map misalignment caused by pose errors by encoding the feature into multiple scales and then aggregating them into a final collaboration feature. In this context, a spatiotemporal awareness CP framework SCOPE [196] was designed to generate robust representations while accounting for localization errors through confidence-aware multi-scale feature interaction. Similarly, the Multi-scale Window Attention Module in V2X-ViT [145] was proposed to capture long-range spatial interactions at multiple scales using a pyramid of resolution windows, allowing for local and global spatial feature interactions and improving detection robustness against positioning errors. In EMIFF [149], the Multi-scale Cross Attention and Camera-aware Channel Masking modules are developed to address the negative effects of pose errors by selectively integrating features, applying spatial offsets, and reweighting features based on camera parameters.
@@
Inter-agent object association aims to mitigate localization errors in pose transformation by matching key points or objects from connected agents and determining the optimal relative pose transformation. To facilitate this, the Context-based Matching (CBM) algorithm [197] generated object clusters based on bounding box headings and established correspondences through iterative pairwise comparison to maximize global consensus. The work in [198] handled spatial differences using a data association and spatial error calibration method based on the Kuhn-Munkres algorithm. However, identifying correspondences can be challenging due to visual folding, non-visible objects, and noisy perception. To address this, [199] framed correspondence identification as a graph-matching problem,
TABLE V: Summary of the Works of Cooperative Information Fusion.
@@
CoCa3D [205] 2023 V2X Camera OPV2V$+$, DAIR-V2X, CoPerception-UAVs$+$ Intermediate Detection $\surd$
@@
CPM-LCF [209] 2021 V2X LiDAR -- Late Detection $\times$
@@
employing a masked GNN to manage non-visible objects, explicitly addressing occlusion and limited FoV challenges.
IV-E Cooperative Information Fusion
The effectiveness of CP crucially depends on the design of an efficient fusion strategy, which aims to consolidate shared information to provide coherent and accurate perception results. The following will examine recent fusion methods employed in CP information. For reference, Table V provides a summary of all discussed fusion models in this subsection.
@@ IV-E1 Traditional Fusion
Primary reduction operators, such as summation, maximum, and average pooling, are widely used in cooperative information fusion for multi-agent perception. For example, the Cooper [97] was developed for CAVs to enhance 3D object detection using raw point cloud data, though it involved high data overhead. F-Cooper [203] extended Cooper to the feature level, balancing bandwidth constraints and detection accuracy by using max pooling for V2V voxel feature fusion. Similarly, PillarGrid [204] applied grid-wise max-pooling to integrate deep features for cooperative 3D object detection. CoCa3D [205] used point-wise maximum fusion to combine BEV features from multiple CAVs, enhancing camera-based detection through cooperative depth estimation. Besides using these commonly accepted reduction operators, dedicated fusion modules were designed to fuse BEV features from CAVs as well. SiCP [206] proposed a novel Dual-Perception Network (DP-Net) to assign different weights for LiDAR point cloud provided by different CAVs by extracting crucial gradient information requiring only a few parameters. CoBEVFusion [207] extended the idea by employing a dedicated fusion module to combine heterogeneous features gathered from the camera and LiDAR on two CAVs.
As the late fusion strategy requires a minimum volume of exchanged information between CAVs [211], several traditional fusion approaches employing the late fusion strategy have been explored in practice. In the context of object detection, non-maximum suppression (NMS) serves as a non-parametric method for filtering overlapping bounding boxes based on confidence scores without requiring training or fitting parameters. V2X-BGN [208] used NMS to aggregate redundant detection results from intelligent agents globally, based on 2D IoU in BEV space. A rule-based fusion scheme for CPMs was proposed in [209] to associate locally detected objects with those identified via V2X communication, designed for scenarios where CAVs generate CPMs based on ETSI standards. This scheme utilized a fusion centre to merge and broadcast CPMs to nearby CAVs for further perception tasks.
The non-parametric fusion methods offer minimal computational costs but often fail to capture data correlations. In contrast, parametric methods generally provide more accurate and reliable fusion by incorporating measurements and their uncertainties. For instance, the Covariance Intersection (CI) algorithm can consistently fuse estimates without prior knowledge of cross-correlations. Building on this concept, [210] introduced a CI-based fusion strategy for CP, achieving consistent fusion of local sensor data and object information from other agents through V2X communication.
@@ IV-E2 Graph-based Fusion
In V2VNet [140], each vehicle used a fully connected GNN [223] as an aggregation module for data fusion from various viewpoints. Messages were aggregated at each node using a mask-aware accumulation operator, considering overlapping FoVs. However, it employed a scalar-valued edge weight to indicate the importance of the neural message from one agent to another, which could not effectively model the strength of collaboration between two agents. On the other hand, DiscoNet [144] employed an edge encoder to establish correlations between feature maps from different agents, generating matrix-valued edge weights. The edge encoder concatenated feature maps and applied convolutional layers for downsampling, followed by a softmax operation for normalization. Knowledge distillation enhanced training by aligning intermediate fusion-based student model features with the early fusion-based teacher model, improving feature abstraction and object detection performance.
The Heterogeneous Multi-agent Self-attention (HMSA) module in V2X-ViT [145] fuse cooperative feature by associating agent types with nodes and edges in a directed graph and computing importance weights between nodes accordingly. To further enhance CP performance, HYDRO-3D [212] extends the V2X-ViT to create a hybrid object detection and tracking CP model by combining V2X-ViT’s detection features with historical tracking data extracted using a spatial-temporal pyramidal 3D network. The combined features improve robustness and enhance detection accuracy, as demonstrated in experiments.
Traditional feature fusion methods, such as concatenation or summation, provide only fixed linear fusion and fail to account for the varying importance of feature maps. To address this, [213] proposed a graph attention feature fusion network that selectively emphasizes important regions within feature maps by employing both channel and spatial attention-based aggregation. This approach adaptively highlights informative areas, enhancing the ego vehicle’s feature representation and improving object detection precision.
@@ IV-E3 Attention-based Fusion
The attention mechanism introduced in the transformer model has found widespread use in determining fusion weights. The Point Transformer [162] effectively learns compact representations from 3D point clouds, excelling at capturing non-local interactions and generating permutation-invariant representations. It is well-suited for aggregating LiDAR point clouds from multiple agents. In COOPERNAUT [146], a point transformer-based representation aggregator was introduced for information fusion, consisting of a voxel max-pooling operation for spatial aggregation and a point transformer block for fusing multi-agent perception data. Similarly, V2VFormer [216] and V2VFormer++ [217] also utilized channel and spatial information based attentions to fuse the features exacted from multiple CAVs. Together with knowledge distillation technology, several attention modules were employed to effectively transfer valuable knowledge from multiple excellent teacher models in different views and then aggregated into the student model on the ego CAV [219, 218].
@@
The aforementioned attention-based models have significantly advanced cooperative information fusion by focusing primarily on spatial dimensions. However, their reliance on single-frame prediction neglects historical context and overlooks important temporal semantic cues, limiting their ability to detect fast-moving objects due to point cloud sparsity. To address these limitations, SCOPE [196] introduced a Context-aware Information Aggregation model that uses a pyramid LSTM to incorporate spatiotemporal data from previous ego agent frames, generating refined context-aware features. These features are fused with cooperative features based on complementary contributions and spatial attention maps obtained through max pooling, which may not fully capture correlations among informative features. To overcome this, IoSP-CP [121] proposed the Historical Prior Hybrid Attention (HPHA) fusion algorithm, comprising a multi-scale transformer and a short-term attention module. The multi-scale transformer extracts attention weights at different spatial scales, while the short-term attention mechanism captures temporal dependencies and identifies discriminative features, enhancing informative features while reducing background noise.
TABLE VI: Summary of the Works on Issues of Cooperative Perception and Corresponding Methods.
@@
UDT2T [226] 2022 V2V -- CARLA Late Tracking $\times$
@@
LC V2VAM$+$LCRN [235] 2022 V2V LiDAR OPV2V Intermediate Detection $\times$
@@
TrustEsti [246] 2019 V2V Camera/LiDAR -- -- -- $\times$
@@
DP FedAD [250] 2023 V2I Camera Udacity -- Turning Angle Prediction $\times$
FedBEVT [251] 2023 V2X Camera CARLA&OpenCDA -- Segmentation $\surd$
FL-CCS [252] 2023 V2X -- -- -- -- $\times$
FedDWA [253] 2024 V2X Camera CARLA&OpenCDA -- Segmentation $\times$
PU CoPEM [254] 2022 V2X -- ViSTA Late Detection $\times$
DMMCP [255] 2023 V2X Camera/LiDAR -- Late Detection, Tracking $\times$

### v5

@@ IV-C1 DSRC
Despite its early adoption, DSRC is increasingly being deprioritized in regulatory and industry roadmaps in favor of C-V2X, which offers greater scalability, network support, and integration with future 5G systems. Notably, the U.S. Federal Communications Commission (FCC) reallocated the 5.9 GHz spectrum in 2024, designating C-V2X as the primary technology for Intelligent Transportation Systems (ITS) and phasing out DSRC-based use in this band [80].
@@ IV-C3 Impact of V2X on CP
CP demands real-time, high-quality data, often involving high-resolution inputs from sensors such as LiDAR and radar, which places significant data rate requirements on communication systems. Achieving high data rates typically requires large bandwidth, stable transmission channels, and precise channel detection and estimation. DSRC operates within the 5.9 GHz frequency band with a total bandwidth of 75 MHz, divided into seven 10 MHz channels (5.8505.9255.850-5.925 GHz). However, due to high vehicle mobility, dynamic environmental changes, and multipath delay spread, DSRC’s communication channels are highly time- and frequency-selective [84]. Studies estimate that up to 50% of the coherence bandwidth is in the order of 1 MHz, with coherence times as short as 0.2 ms [84], limiting reliable transmission times and increasing dropout rates. Similar constraints affect C-V2X, which operates on a narrower 30 MHz spectrum (5.8955.9255.895-5.925 GHz) [85]. The PC5 mode of C-V2X offers an advantage by utilizing a blind hybrid automatic repeat request (HARQ) technique that combines turbo coding (providing better performance than the convolutional coding used in DSRC) with automatic repeat request error control and soft-combining [86].
Addressing these challenges requires improved network planning and advanced channel detection and estimation techniques. For DSRC, an analytical model was proposed in [87] to assess key metrics such as Channel Busy Ratio and Packet Delivery Ratio over distance, facilitating optimized network planning. Additionally, enhancements to IEEE 802.11p were made in [88] through an OFDM-IM system in which pilot signals were strategically positioned and inactive channels were leveraged to increase detection accuracy. In [89], a low-complexity, high-speed deep neural network-augmented least square algorithm was introduced, resulting in reduced model size and latency, as well as improved accuracy under varied noise levels and channel conditions. For C-V2X, a balanced resource allocation strategy was proposed in [90], in which resources were evenly distributed across subframes, leading to decreased packet loss and enhanced data throughput. Similarly, a pilot design incorporating zero-correlation zone sequences and a time-domain channel estimation approach was suggested in [91] to mitigate co-channel interference in C-V2X systems.
@@
To address congestion in DSRC, adjustments to vehicle settings have been proposed in [93] based on information from supplementary channels, enabling lead vehicles to monitor hidden nodes within a group. Furthermore, a congestion-aware forwarding scheme was presented in [94] to optimize multi-hop forwarding reliability and reduce latency under heavy network load. In C-V2X, a rate control mechanism has been introduced in [91], allowing power control to actively contribute to congestion management, while a DQN-based adaptive congestion control algorithm was suggested in [95], in which a central controller is used to train and distribute the policy, thereby improving QoS compliance over ETSI’s Distributed Congestion Control in dense traffic scenarios. For synchronization, a decentralized clock synchronization protocol was introduced in [96] to enhance DSRC’s resilience against GPS spoofing and to improve reliability in the absence of GNSS signals. Meanwhile, a synchronization signal design using generalized reciprocal sequences for mmWave C-V2X was proposed in [97]. This approach is shown to enhance OFDM-based V2V synchronization by achieving high detection accuracy, minimizing interference, and reducing computational complexity.
@@ IV-D Cooperative Information Alignment
Cooperative information alignment is crucial for CP systems in autonomous driving to ensure that shared data from different sources is accurately represented within a common coordinate system [98]. This process is essential for fusing information from multiple agents to achieve a unified understanding of the environment.
@@ IV-D2 Alignment for Pose Errors
TABLE III: Summary of the Works of Cooperative Information Alignment for Pose Errors.
@@
DMGM [111] 2023 V2V Camera CARLA & SUMO Early - $\surd$
@@
Many efforts have been undertaken to address pose errors, as summarized in Table III. A widely adopted approach involves statistical learning models. For instance, the method presented in [100] corrected relative locations between two CAVs by matching key points in their perception results with the Random Sample Consensus (RANSAC) algorithm [112]. While effective, this approach required the setting of problem-specific thresholds. Based on V2VNet [54], a consistency module [101] was introduced to enhance relative pose estimations by structuring the global coherence of absolute poses as a Markov random field and incorporating Bayesian reweighting. Unfortunately, this model relies on ground-truth poses during the training phase, which may be limited in practical scenarios. In CoAlign [102], an Agent-Object Pose Graph Optimization method was proposed to improve pose consistency through pose graph construction and optimization techniques. Similarly, in RoCo [103], the pose correction problem was modeled as an object-matching task, wherein common objects detected by different agents were reliably associated, and agent poses were iteratively adjusted to minimize alignment errors. Both CoAlign and RoCo effectively mitigated relative pose errors without requiring precise supervision during the training process.
@@
Point cloud registration methods typically focus on refining the Iterative Closest Point (ICP) algorithm [113], widely used for aligning 3D point clouds. In [104], an ICP-based matching algorithm aligned two point sets, one representing 3D bounding boxes detected by the Ego and the other by the sender, with the relative transformation formulated as an optimal transport problem [114] to minimize transportation costs between source and target points. In V2X-PC [50], point clusters from spatial and temporal dimensions were aligned, and parameter-free solutions were proposed to adapt to varying noise levels without requiring finetuning. In CoBEVGlue [105], a spatial alignment module named BEVGlue was introduced to estimate relative poses between agents by matching co-visible objects without relying on an external localization system. In contrast to traditional point cloud registration methods, BEVGlue required only the transmission of bounding boxes and tracking identifiers, thereby reducing communication bandwidth while maintaining high-quality matching through verification of spatial relationships between matched nodes.
@@
High-definition maps represent a valuable resource for achieving precise spatial alignment, primarily due to their ability to provide accurate self-vehicle localization. In this regard, a map-based CP framework called the “map container” was introduced in [106], which establishes benchmarks for spatiotemporal transformations by automatically mapping multi-agent perceptual information into the map coordinate space using a map-matching algorithm. However, the implementation of this algorithm was dependent on monocular cameras and GNSS/IMU devices. In contrast, FreeAlign [107] was proposed to estimate relative poses by leveraging the invariant geometric structure formed by commonly perceived and observed objects across agents, thereby avoiding reliance on expensive high-end global localization systems.
@@
Multi-scale feature interaction has been shown to effectively mitigate feature map misalignment caused by pose errors by encoding features at multiple scales and subsequently aggregating them into a final collaborative representation. In this context, a spatiotemporal awareness CP framework, SCOPE [108], was designed to generate robust representations while accounting for localization errors through confidence-aware multi-scale feature interaction. Similarly, in V2X-ViT [59], a Multi-scale Window Attention Module was proposed to capture long-range spatial interactions at multiple scales through a pyramid of resolution windows, enabling both local and global spatial feature interactions and improving detection robustness against positioning errors. In EMIFF [63], the Multi-scale Cross Attention and Camera-aware Channel Masking modules were developed to counter the negative effects of pose errors by selectively integrating features, applying spatial offsets, and reweighting features based on camera parameters.
@@
Inter-agent object association is aimed at mitigating localization errors in pose transformation by matching key points or objects from connected agents and determining the optimal relative pose transformation. To facilitate this, the Context-based Matching (CBM) algorithm [109] was developed, in which object clusters were generated based on bounding box headings and correspondences were established through iterative pairwise comparison to maximize global consensus. Spatial differences were addressed in [110] using a data association and spatial error calibration method based on the Kuhn-Munkres algorithm. However, the identification of correspondences remains challenging due to visual folding, non-visible objects, and noisy perception. To overcome these issues, correspondence identification was framed as a graph-matching problem in [111], and a masked GNN was employed to handle non-visible objects, explicitly accounting for occlusion and limited field-of-view challenges.
IV-E Cooperative Information Fusion
The effectiveness of CP crucially depends on the design of an efficient fusion strategy, which aims to consolidate shared information to provide coherent and accurate perception results. The following will examine recent fusion methods employed in CP information. For reference, Table IV provides a summary of all discussed fusion models in this subsection.
TABLE IV: Summary of the Works of Cooperative Information Fusion.
@@ IV-E Cooperative Information Fusion
CoCa3D [118] 2023 V2X Camera OPV2V++, DAIR-V2X, CoPerception-UAVs++ Intermediate Detection $\surd$
@@
CPM-LCF [122] 2021 V2X LiDAR - Late Detection $\times$
@@ IV-E1 Traditional Fusion
Primary reduction operators, such as summation, maximum, and average pooling, are widely used in cooperative information fusion for multi-agent perception. For example, the Cooper [115] was developed for CAVs to enhance 3D object detection using raw point cloud data, though it involved high data overhead. This was extended in F-Cooper [116], where feature-level fusion was employed to balance bandwidth constraints and detection accuracy through the use of max pooling for vehicle-to-vehicle voxel feature fusion. Similarly, grid-wise max pooling was applied in PillarGrid [117] to integrate deep features for cooperative three-dimensional object detection. In CoCa3D [118], point-wise maximum fusion was utilized to combine bird’s-eye view (BEV) features from multiple CAVs, thereby enhancing camera-based detection via cooperative depth estimation. In addition to these commonly adopted reduction operators, dedicated fusion modules have also been designed for fusing BEV features across CAVs. In SiCP [119], a Dual-Perception Network (DP-Net) was proposed to assign different weights to LiDAR point clouds from different CAVs by extracting crucial gradient information, requiring only a small number of parameters. This concept was extended in CoBEVFusion [120], where a dedicated fusion module was employed to integrate heterogeneous features collected from the camera and LiDAR sensors on two CAVs.
As the late fusion strategy requires a minimum volume of exchanged information between CAVs, several traditional fusion approaches employing the late fusion strategy have been explored in practice. In the context of object detection, non-maximum suppression (NMS) serves as a non-parametric method for filtering overlapping bounding boxes based on confidence scores without requiring training or fitting parameters. In V2X-BGN [121], NMS was applied to aggregate redundant detection results from intelligent agents globally, based on 2D IoU in BEV space. A rule-based fusion scheme for cooperative perception messages (CPMs) was proposed in [122] to associate locally detected objects with those identified via vehicle-to-everything communication. This scheme was designed for scenarios in which CAVs generate CPMs according to the European Telecommunications Standards Institute (ETSI) specifications and employ a fusion centre to merge and broadcast the CPMs to nearby CAVs for subsequent perception tasks.
The non-parametric fusion methods offer minimal computational costs but often fail to capture data correlations. In contrast, parametric methods generally provide more accurate and reliable fusion by incorporating measurements and their uncertainties. For instance, the Covariance Intersection (CI) algorithm can consistently fuse estimates without prior knowledge of cross-correlations. Building on this concept, a CI-based fusion strategy for cooperative perception was introduced in [123], in which local sensor data and object information received from other agents via V2X communication were consistently fused.
@@ IV-E2 Graph-based Fusion
In V2VNet [54], each vehicle used a fully connected GNN [136] as an aggregation module for data fusion from various viewpoints. Messages were aggregated at each node using a mask-aware accumulation operator, considering overlapping Field of Views (FoVs). However, it employed a scalar-valued edge weight to indicate the importance of the neural message from one agent to another, which could not effectively model the strength of collaboration between two agents. In contrast, in DiscoNet [58], an edge encoder was employed to establish correlations between feature maps from different agents by generating matrix-valued edge weights. Feature maps were concatenated and processed through convolutional layers for downsampling, followed by a softmax operation for normalization. Additionally, knowledge distillation was used to enhance training by aligning the intermediate features of a fusion-based student model with those of an early fusion-based teacher model, thereby improving feature abstraction and object detection performance.
The Heterogeneous Multi-agent Self-attention (HMSA) module in V2X-ViT [59] was designed to fuse cooperative features by associating agent types with nodes and edges in a directed graph, and by computing importance weights between nodes accordingly. To further enhance cooperative perception performance, HYDRO-3D [125] was developed as an extension of V2X-ViT to construct a hybrid object detection and tracking model. In this framework, detection features from V2X-ViT were combined with historical tracking data extracted using a spatial-temporal pyramidal three-dimensional network. The fused features were shown to improve robustness and enhance detection accuracy, as demonstrated through experimental evaluation.
Traditional feature fusion methods, such as concatenation and summation, offer only fixed linear fusion and do not account for the varying importance of feature maps. To address this limitation, a graph attention feature fusion network was proposed in [126], in which important regions within feature maps were selectively emphasized through channel and spatial attention based aggregation. Through this approach, informative areas were adaptively highlighted, leading to enhanced feature representation for the ego vehicle and improved object detection precision.
@@ IV-E3 Attention-based Fusion
The attention mechanism introduced in the transformer model has been widely adopted for determining fusion weights. The Point Transformer [73] was shown to effectively learn compact representations from three-dimensional point clouds, demonstrating strong capability in capturing non-local interactions and producing permutation-invariant representations. This made it well-suited for aggregating LiDAR point clouds from multiple agents. In COOPERNAUT [60], a point transformer-based representation aggregator was introduced for information fusion, consisting of a voxel max-pooling operation for spatial aggregation and a point transformer block for fusing multi-agent perception data. Similarly, in V2VFormer [129] and V2VFormer++ [130], attentions based on channel and spatial information were employed to fuse features extracted from multiple CAVs. Together with knowledge distillation technology, several attention modules were employed to effectively transfer valuable knowledge from multiple excellent teacher models in different views and then aggregated into the student model on the ego CAV [132, 131].
@@
The aforementioned attention-based models have significantly advanced cooperative information fusion by focusing primarily on spatial dimensions. However, their reliance on single-frame prediction neglects historical context and overlooks important temporal semantic cues, limiting their ability to detect fast-moving objects due to point cloud sparsity. To address these limitations, a Context-aware Information Aggregation model was introduced in SCOPE [108], in which a pyramid LSTM network was employed to incorporate spatiotemporal data from previous frames of the ego agent, generating refined context-aware features. These features were fused with cooperative features based on complementary contributions and spatial attention maps obtained through max pooling, which may not fully capture correlations among informative features. To overcome this shortcoming, the Historical Prior Hybrid Attention (HPHA) fusion algorithm was proposed in IoSP-CP [34], comprising a multi-scale transformer and a short-term attention module. In this framework, attention weights were extracted at different spatial scales using the multi-scale transformer, while the short-term attention module was used to capture temporal dependencies and identify discriminative features, thereby enhancing informative features and suppressing background noise.
TABLE V: Summary of the Works on Issues of Cooperative Perception and Corresponding Methods.
@@
UDT2T [139] 2022 V2V - CARLA Late Tracking $\times$
@@
LC V2VAM++LCRN [148] 2022 V2V LiDAR OPV2V Intermediate Detection $\times$
@@
TrustEsti [159] 2019 V2V Camera/LiDAR - - - $\times$
@@
DP FedAD [163] 2023 V2I Camera Udacity - Turning Angle Prediction $\times$
FedBEVT [164] 2023 V2X Camera CARLA&OpenCDA - Segmentation $\surd$
FL-CCS [165] 2023 V2X - - - - $\times$
FedDWA [166] 2024 V2X Camera CARLA&OpenCDA - Segmentation $\times$
PU CoPEM [167] 2022 V2X - ViSTA Late Detection $\times$
DMMCP [168] 2023 V2X Camera/LiDAR - Late Detection, Tracking $\times$